### 10 Feb 2024 v2

# Making Language Models Better Tool Learners with Execution Feedback

Abstract
Tools serve as pivotal interfaces that enable humans to understand and reshape the environment. With the advent of foundation models, AI systems can utilize tools to expand their capabilities and interact with the real world. Existing tool learning methodologies, encompassing supervised fine-tuning and prompt engineering approaches, often induce large language models to utilize tools indiscriminately, as complex tasks often exceed their own competencies. However, introducing tools for simple tasks, which the models themselves can readily resolve, can inadvertently propagate errors rather than enhance performance. This leads to the research question: can we teach language models when and how to use tools? To meet this need, we propose Tool leaRning wIth exeCution fEedback (Trice), a two-stage end-to-end framework that enables the model to continually learn through feedback derived from tool execution, thereby learning when and how to use tools effectively. Experimental results, backed by further analysis, show that Trice can make the large language model selectively use tools by improving the accuracy of tool usage while enhancing insufficient tool learning and mitigating excessive reliance on tools [^1] Code and datasets are in https://github.com/zjunlp/TRICE..
Making Language Models Better Tool Learners with Execution Feedback
Shuofei ${}^{\spadesuit}$, Honghao ${}^{\spadesuit}$, Chengfei ${}^{\clubsuit}$, Qianghuai ${}^{\clubsuit}$, Huajun ${}^{\spadesuit\heartsuit}$, Ningyu ${}^{\spadesuit}$thanks:     Corresponding Author. ${}^{\spadesuit}$ Zhejiang University ${}^{\heartsuit}$ Donghai Laboratory ${}^{\clubsuit}$ Alibaba Group {shuofei,guihonghao,huajunsir,zhangningyu}@zju.edu.cn
1 Introduction
The recent rapid advancement of foundation models (Brown et al., 2020; Ouyang et al., 2022; Chowdhery et al., 2022; Qiao et al., 2023; Zhao et al., 2023b) makes it practical for AI machines to create (Cai et al., 2023; Qian et al., 2023) and utilize tools effectively (Shen et al., 2023; Lu et al., 2023), which greatly transcends their inherent limitations in various underlying areas, including arithmetic (Cobbe et al., 2021; Parisi et al., 2022), knowledge updating (Sun et al., 2023; Zhao et al., 2023a), multi-modal semantic analysis (Wu et al., 2023; Driess et al., 2023), etc.
Figure 1: Large language model learns to use tools from execution feedback.
Existing research has shed light on the potential of Large Language Models (LLMs) to exhibit a promising level of dexterity and finesse in tool use (Qin et al., 2023a; Wang et al., 2023). Prior works view tools as external resources to augment LLMs for better performance (Schick et al., 2023; Hao et al., 2023; Patil et al., 2023; Tang et al., 2023) or employ LLMs as a hub for human-tool interaction, responsible for orchestrating the deployment and usage of tools (Shen et al., 2023; Ge et al., 2023; Lu et al., 2023; Driess et al., 2023).
Method LM Model Scale Mechanism Feedback Peft Teacher Unseen
  Toolformer Schick et al. (2023) GPT-J 6B instruct-tuning
ToolkenGPT Hao et al. (2023) LLaMA 13B, 30B fine-tuning
HuggingGPT Shen et al. (2023) ChatGPT ¿=100B prompt
Chameleon Lu et al. (2023) GPT-4 ¿=100B prompt
ChatCoT Chen et al. (2023) ChatGPT ¿=100B prompt
Gorilla Patil et al. (2023) LLaMA 7B instruct-tuning
ToolLLaMA Qin et al. (2023b) LLaMA 7B instruct-tuning
GPT4Tools Yang et al. (2023) Vicuna 13B instruct-tuning
Trice (ours) ChatGLM Alpaca Vicuna 6B, 7B instruct-tuning reinforcement learning
Table 1: Comparison of related works. Mechanism denotes how the LM learns to invoke tools. Feedback stands for whether the LM learns from execution feedback. Peft means the parameter efficient tuning. Teacher expresses whether learning from a powerful teacher like ChatGPT. Unseen indicates the zero-shot capability on unseen tools.
Despite the empirical success of previous work, a critical issue remains: LLMs often do not understand when and how to properly use which tools. On one hand, the use of tools is necessary to augment LLMs when facing complex problems that surpass their inherent capabilities. On the other hand, for simpler problems that can readily be solved by the models themselves, introducing tools can paradoxically propagate errors rather than enhance performance. These errors can include but are not limited to, improper selection of tool types, generation of incorrect tool inputs, and ineffective utilization of tool return results. Intuitively, it’s crucial for LLMs to develop an awareness of when tools are necessary and when they are not, and to be able to make decisions about selecting the most appropriate tools for the task at hand.
To address the above issues, we propose Tool leaRning wIth exeCution fEedback (Trice) as shown in Figure 1, a two-stage end-to-end framework that enables the model to continually learn through feedback derived from execution, thereby learning when and how to use tools effectively. Specifically, we first prepare a dataset that helps discern when tool usage is necessary for LLMs and when it is not. Given the lack of gold labels, we utilize ChatGPT (OpenAI, 2022) to automatically generate tool usage APIs. Then, we introduce a two-stage training strategy to teach the model when to use tools: 1) Behavior Cloning. We conduct instruct-tuning on the dataset to let the model imitate the tool-using behavior. 2) Reinforcement Learning with Execution Feedback (RLEF). We further reinforce the model with execution feedback by aligning it with desirable candidate responses, guiding the model to selectively use tools to avoid error propagation. We detail the main difference of Trice with related works in Table 1.
We train and evaluate Trice on various tasks and backbone models. Experimental results and further analyses demonstrate that Trice successfully instructs the model to judiciously use tools, simultaneously enhancing insufficient tool learning, reducing excessive reliance on tools, and improving the accuracy of tool usage. In summary, the key contributions of our study are as follows:
  • We introduce Trice, a two-stage end-to-end training framework that leverages execution feedback to help LLMs become more proficient tool learners.
  • We perform superior on eight benchmark datasets of four tasks with various models.
  • Extensive empirical analysis demonstrates that Trice can guide the model in judicious tool use, thereby enhancing insufficient tool use, reducing excessive dependency on tools, and improving the effectiveness of tool use.
2 Related Work
Tool Learning.
Though possessing remarkable capabilities (Qiao et al., 2023; Yao et al., 2023), LLMs still struggle in many basic aspects where much smaller and simpler tools may precisely excel. Under this circumstance, a new paradigm, called Tool Learning (Qin et al., 2023a), is born to combine the strengths of both LLMs and specialized tools. Some works (Driess et al., 2023; Shen et al., 2023; Lu et al., 2023) regard LLMs as a decision-making hub for compositional tool using which can be called Tool-Oriented Learning (Qin et al., 2023a), while others (Gao et al., 2022; Liu et al., 2023; Schick et al., 2023) treat tools as complementary resources to extend the power of LLMs which can be called Tool-Augmented Learning (Mialon et al., 2023; Qin et al., 2023a). Despite their success, tool-augmented approaches tend to force LLMs to use tools mindlessly regardless of whether they actually need tools for help. This may, in some scenarios, steer LMs to erroneously choose the type of tools or the way to use tools, making the loss outweighs the gain. Compared to previous works, we try to make LMs better tool learners by teaching them to use tools selectively instead of blindly.
Figure 2: The overview of our proposed framework Trice. In stage-I (Behavior Cloning), We conduct instruct-tuning on the dataset to let the model imitate the tool-using behavior. In stage-II (RLEF), we further reinforce the model with tool execution feedback by aligning it with desirable candidate responses.
Learning from Feedback.
An intuitive approach of tool learning is to fit LLMs on examples with human-labeled tools directly (Torabi et al., 2018; Li et al., 2022). However, this is often impractical to annotate every possible scenario (Codevilla et al., 2019) and difficult to generalize to new ones. It is worth noting that humans generally have the ability to correct and reflect on their own behavior from trial and error (Allen et al., 2019). Intuitively, feedbacks from the environments or humans enable LLMs to understand the impact of their actions and adapt their behavior accordingly. Reinforcement learning (RL) excels at enabling models to learn decision-making abilities in complex environments through feedback (Schrittwieser et al., 2020; Yao et al., 2022; Ge et al., 2023). Ouyang et al. (2022) apply a state-of-the-art RL algorithm, PPO (Schulman et al., 2017), to align LLMs with human feedback. Liu et al. (2022) reinforce knowledge for commonsense question answering with a fixed QA model providing feedback. Yuan et al. (2023); Lee et al. (2023) offer a promising alternative that leverages powerful off-the-shelf LLMs to generate preferences. Compared to the previous feedback framework, we introduce RLEF for tool learning which reinforces the LLMs with the execution result of tools.
3 Methodology
Problem Overview.
We mainly focus on four kinds of tasks, with each instance in the format of $x=(s,q,t,a)$, where $s$ denotes the specialized instruction of each task, $q$ refers to the question, $t$ stands for the tool API and $a$ is the gold answer. Following an instruction-following paradigm, the complete input of the LLM is as follows:
$\mathtt{input}=[s,q],$ (1)
where [,] stands for text concatenating. In terms of the output, when LLM deems that no tool is necessary, it generates the answer $a$. Conversely, if the model identifies the need for a tool, it produces the tool API $t$, which encompasses the specific type of tool and its corresponding input:
$\mathtt{output=}\begin{cases}a&use\_tool=\mathtt{false}\\ t&use\_tool=\mathtt{true}\end{cases}$ (2)
The detailed format of each task is shown in A.1.
Given the problem, the main challenges lie in 1) determining the LLM when to or not to harness tools for help and 2) how to impart the ability to the LLM to make selective use of tools. For the former, we allow the untrained model to directly infer answers, considering the correct ones as not requiring tools and the incorrect ones as indicating the need for tool assistance. For the latter, we adopt Trice, a two-stage training strategy. In the first stage, we use Behavior Cloning to make the model imitate tools invoking. Building upon this, we continue to train the model for selective tool usage with RLEF in the second stage. The overview of our method is illustrated in Figure 2. Please note that all symbols are globally defined in sections 3&4.
Data Preparation.
The data preparation follows the principles outlined in Eq.1&2. Given a raw initial dataset $\mathcal{D}_{\textrm{init}}=\{(q,a)\}_{i=1}^{|\mathcal{D}_{\textrm{init}}|}$ from the benchmark, we utilize LLM without fine-tuning to generate predictions. Since we do not have gold labels for tool APIs, we employ ChatGPT (OpenAI, 2022) to generate pseudo-labels under few-shot prompting. Specifically, we generate tool API labels $t=\mathtt{tool_{name}(tool_{input})}$ for questions where the generated predictions are incorrect. For questions with correct predictions, we directly set $t=\mathtt{None}$ to indicate that tool APIs are not required. We design particular instructions $s$ tailored to each task. In the end, we obtain $\mathcal{D}_{\textrm{tool}}=\{(s,q,t,a)\}_{i=1}^{|\mathcal{D}_{\textrm{tool}}|}$ according with Eq.1&2 containing the tool demand information of the specific LLM as we desire [^2] For more details of data preparation, please refer to A.2..
Training.
As shown in Figure 2, based on $\mathcal{D}_{\textrm{tool}}$, we conduct a two-stage training approach: I) Behavior Cloning3.1). In this stage, we teach the model to imitate the tool usage behavior by fine-tuning it on $\mathcal{D}_{\textrm{tool}}$ in an instruct-tuning manner. This empowers the model with preliminary functionality of tool API invocation. II) Reinforcement Learning with Execution Feedback3.2). Drawing inspiration from fine-tuning with human feedback (Ouyang et al., 2022), we continue to reinforce our model obtained in stage I with execution feedback by steering it to align with desirable candidate responses.
3.1 Training Stage I: Behavior Cloning
During the behavior cloning stage, we aim to enable the LLM to master the schema of tool API invocation and develop preliminary skills in selectively utilizing tools. We perform supervised fine-tuning on $\mathcal{D}_{\textrm{tool}}$ in this stage.
Specifically, for the model $p_{\mathrm{LM}}$ with tunable parameters $\theta$, the training loss of stage I can be formulated as:
$\mathcal{L}_{\textrm{clone}}(\theta)=\sum_{(s,q,t,a)\in\mathcal{D}_{\textrm{tool}}}-\log p_{\mathrm{LM}}(o|s,q;\theta),$ (3)
where $o$ is the specified output of the model as defined in Eq.2. The final parameterized model of this stage is denoted as $\theta_{\textrm{clone}}$.
3.2 Training Stage II: RLEF
In stage II, we continue to optimize $\theta_{\textrm{clone}}$ with execution feedback, so as to enhance its capability to selectively utilize tools and improve the accuracy of decision-making regarding tool types and corresponding inputs.
Overall Loss.
Following Yuan et al. (2023), for each question $q$, we have $k$ different candidate responses $y_{i}$ $(1\leq i\leq k)$ marshaled from other LLMs (e.g. ChatGPT, LLaMA) or human experts. We apply a reward strategy to score each $y_{i}$ with $r_{i}=R(a,y_{i})$ where $a$ is the gold answer of question $q$. Our goal is to instruct the LLM to determine the more desirable response by aligning with scores $\{r_{i}\}_{k}$. So we then score each $y_{i}$ with the LLM:
$p_{i}=\frac{\sum_{t}\log p_{\mathrm{LM}}(y_{i,m}|q,y_{i,<m};\theta_{\textrm{clone}})}{||y_{i}||},$ (4)
where $m$ denotes the $m$th token of $y_{i}$, $p_{i}$ is the conditional log probability of $y_{i}$ and $||y_{i}||$ refers to the length-normalized factor.
To facilitate the LLM in learning the correct score ordering of different $y_{i}$, we introduce a ranking loss during training:
$\mathcal{L}_{\textrm{rank}}=\sum_{r_{i}<r_{j}}\max(0,p_{i}-p_{j}).$ (5)
Meanwhile, in order to prevent the model from deviating too far from the original parameters and generating unreasonable tool API invocation structure, we reintroduce the supervised fine-tuning loss:
$\mathcal{L}_{\textrm{sft}}=-\sum_{m}\log p_{\mathrm{LM}}(o_{m}|s,q,o_{<m};\theta_{\textrm{clone}}).$ (6)
Finally, the overall RLEF loss is defined as follows:
$\mathcal{L}_{\textrm{RLEF}}=\alpha\cdot\mathcal{L}_{\textrm{rank}}+\mathcal{L}_{\textrm{sft}},$ (7)
where $\alpha$ is a hyperparameter that determines the proportion of the rank loss.
Reward Strategy.
The reward strategy aims to give each $y_{i}$ an $r_{i}$ and rank them accordingly for a given question $q$. We view the output $o$ regulated in $\mathcal{D}_{\textrm{tool}}$ as the pseudo-human-expert (gold) response. Then the reward strategy is derived from two indicators: 1) accuracy of the answer and 2) consistency of tool usage with the gold response. Specifically, we employ a five-level scoring strategy. We assign the gold response with the maximum score. For the remaining, assuming that the correctness of the response is denoted as $\mathtt{True}$ for correct answers and $\mathtt{False}$ for incorrect answers, and whether the use of tools aligns with the gold response is denoted as $\mathtt{Yes}$ for alignment and $\mathtt{No}$ for misalignment, to ensure accurate and selective tool usage, our scoring is prioritized as follows:
$\mathtt{True}\mathtt{Yes}>\mathtt{True}\mathtt{No}>\mathtt{False}\mathtt{Yes}>\mathtt{False}\mathtt{No}.$
If two responses share the same state, they would receive the same score.
Task Tool Datasets
Math Reasoning Calculator ASDiv (Miao et al., 2020) SVAMP (Patel et al., 2021) GSM8K (Cobbe et al., 2021)
Question Answering WikiSearch WebQuestions (Berant et al., 2013) NaturalQuestions (Kwiatkowski et al., 2019) TriviaQA (Joshi et al., 2017)
LAMA QA Model T-REx (Petroni et al., 2019)
Multilingual QA Translator MLQA (Lewis et al., 2020)
Table 2: Tasks, datasets and the corresponding tools.
4 Experiments
4.1 Experimental Settings
Tasks.
As shown in Table 2, we mainly evaluate our method on four tasks with each task specified to an external tool. Due to limited computational resources, we randomly sample train and test sets from each dataset to reduce the data scale. We display the detailed data distribution for each task in Figure 11. Following Schick et al. (2023), we use a more lenient evaluation criterion than exact match. We simply check for the last number predicted by the model for the math reasoning task and check whether the correct answer is within the first twenty words for other tasks.
Candidate Response Generation.
We collect five responses for each question from four different models, e.g. ChatGPT, InstuctGPT, Vicuna-7B, Alpaca-7B, and the output regulated in $\mathcal{D}_{\textrm{tool}}$ as the gold response. To differentiate whether or not to use tools among candidate responses, we compel ChatGPT and InstructGPT to utilize tools while allowing Alpaca and Vicuna to make the choice of using tools. For ChatGPT and InstructGPT, we prompt them with instructions and few-shot examples, and for Alpaca-7B and Vicuna-7B, we fine-tune them on $\mathcal{D}_{\textrm{tool}}$ with LoRA (Hu et al., 2022) for a few steps in order to equip them with initial abilities for question answering and tool generation [^3] For more details of candidate response generation, please refer to A.3..
Setting Model Math Reasoning Question Answering LAMA Multilingual QA Avg.
ASDiv SVAMP GSM8K WebQ NaturalQ TriviaQA T-REx MLQA
  GPT-3.5 64.6 62.0 19.8 46.4 15.0 41.3 58.7 34.4 42.8
  ChatGLM (Zero-Shot) 30.8 30.5 6.3 12.1 1.6 3.9 21.8 36.5 17.9
Prompt Based ChatGLM (Few-Shot) 34.5 30.5 7.1 11.9 1.9 3.5 23.5 36.7 18.7
ChatGLM (0% Tool) 44.2 35.5 7.2 14.9 9.5 11.2 30.6 37.7 23.9
Supervised Fine-Tuning ChatGLM (100% Tool) 68.2 59.5 11.8 12.5 9.9 13.8 26.8 35.9 29.8
ChatGLM (Trice-split) 72.9 64.0 12.4 15.2 11.6 15.2 32.7 37.3 32.7
Trice Based ChatGLM (Trice-mix) 75.6 65.5 15.8 18.5 13.7 29.0 34.7 41.7 36.8
  Alpaca (Zero-Shot) 31.2 22.0 3.5 32.8 5.3 15.0 39.7 37.7 23.4
Prompt Based Alpaca (Few-Shot) 38.3 23.5 4.3 33.9 6.0 16.6 41.1 45.5 26.2
Alpaca (0% Tool) 44.0 23.0 5.8 37.6 10.3 20.4 53.1 48.9 30.4
Supervised Fine-Tuning Alpaca (100% Tool) 68.6 44.5 15.6 35.9 16.4 32.6 41.7 46.6 37.7
Alpaca (Trice-split) 73.4 45.0 16.3 38.2 18.6 37.8 54.6 48.2 41.5
Trice Based Alpaca (Trice-mix) 75.2 58.0 21.5 41.4 20.7 41.4 55.2 52.0 45.7
  Vicuna (Zero-Shot) 50.4 33.0 6.4 34.9 7.7 16.7 42.5 35.9 28.4
Prompt Based Vicuna (Few-Shot) 56.1 35.5 6.9 36.2 8.8 17.6 44.2 38.5 30.5
Vicuna (0% Tool) 52.3 38.5 8.1 38.8 11.5 20.8 52.9 44.3 33.4
Supervised Fine-Tuning Vicuna (100% Tool) 69.4 48.0 15.8 37.1 17.5 33.9 45.7 42.1 38.7
Vicuna (Trice-split) 72.6 49.0 16.6 43.2 20.7 40.8 54.1 42.6 42.4
Trice Based Vicuna (Trice-mix) 81.2 60.5 21.8 44.1 21.2 41.6 55.4 49.7 46.9
Table 3: Performance of Trice across various tasks with different backbone models. Zero-Shot: models are directly evaluated. Few-Shot: models are prompted by 3 examples during evaluation. 0% Tool: models are trained purely on question-answer paired data. During the above settings, the model does not rely on tools. 100% Tool: models are trained purely on question-tool paired data. Trice-split: models are trained with Trice separately for each task. Trice-mix: models are trained with Trice by combining training data from all tasks.
Baselines.
We mainly experiment with the following LLMs: 1) GPT-3.5 (OpenAI, 2022). We utilize the text-davinci-003 version of GPT series. 2) ChatGLM-6B (Du et al., 2022), a general language model pre-trained with an autoregressive blank-filling objective. 3) Alpaca-7B (Taori et al., 2023), a model further trained on LLaMA-7B (Touvron et al., 2023) with self-instruction. 4) Vicuna-7B (Chiang et al., 2023), an open-source chatbot trained by fine-tuning LLaMA-7B. For GPT-3.5, we directly utilize the OpenAI API, while for other models, we train them all with LoRA (Hu et al., 2022) for efficiency in both stage-I&II.
[…]
Setups.
All the models are trained for 5 epochs in stage-I and 2 epochs in stage-II. We use the learning rates of {2e-5, 1e-4, 3e-4} for ChatGLM-6B and {2e-5, 1e-4} for Alpaca-7B and Vicuna-7B. The $\alpha$ is set to {0.01, 0.1, 1} for all the models. The detailed hyper-parameters we use are shown in A.4. Since sampling responses and training are separated, our whole training procedure only needs to load one model. All our training can be completed on one 80G A800 GPU within 10 hours.
4.2 Main Results
Selective Tool Learning of Single Tool.
Within each task, we train the model to learn the corresponding tool as shown in Table 2, thereby evaluating the model’s proficiency in handling a single tool. From the rows labeled Trice-split in Table 3, it is evident that training by Trice, Alpaca and Vicuna perform on par with GPT-3.5, exhibiting only a slight decrease of $\downarrow$1.3and $\downarrow$0.4on average. Meanwhile, across all backbone models, Trice-split demonstrates significant improvements compared to the prompt-based baselines, surpassing the few-shot setting with $\uparrow$14.0for ChatGLM, $\uparrow$15.3for Alpaca, and $\uparrow$11.9for Vicuna. This indicates that Trice consistently empowers LLMs to use tools effectively, irrespective of the model architecture and scale. Moreover, whether it is completely independent (0% Tool) or dependent (100% Tool) on tools, supervised fine-tuning fails to beat Trice-based training, which highlights the necessity and efficacy of judicious tool learning.
Selective Tool Learning of Multi-Tools.
Across all tasks, we train the model to simultaneously learn all the tools, assessing its capabilities in multi-tool learning. As indicated in the rows labeled Trice-mix in Table 3, training across tasks achieves state-of-the-art performance by further exceeding the Trice-split with over $\uparrow$4.0average score gains across different models. Meanwhile, both Alpaca and Vicuna outperform GPT-3.5, exhibiting improvements of $\uparrow$2.9and $\uparrow$4.1%, respectively. These results declare the potential of Trice in selective multi-tool learning, which paves the way for expanding the capabilities of LLMs to wisely handle more complex and diverse types of tools.
Model Unseen Dataset Unseen Tool
Calculator QA Model Retriever
MultiArith AddSub SQuAD HotpotQA
  GPT-3.5 51.1 59.5 45.2 36.7
  Vicuna (Zero-Shot) 42.3 44.1 28.6 19.7
Vicuna (Few-Shot) 45.5 49.1 31.2 20.6
Vicuna (Trice-split) 63.1 75.2 30.9
Vicuna (Trice-mix) 66.6 80.5 35.7 27.3
Table 4: Performance to unseen datasets and tools.
Figure 3: Performance of Trice across all tasks at different training stages. Trice-I: only train by Behavior Cloning (instruct-tuning) stage. Trice-II: only train by RLEF (reinforcement learning with execution feedback) stage. Trice-all: train by both Trice-I and Trice-II.
Figure 4: Comparison of tool use rate statistics among different training stages. In the Zero-Shot stage, we consider a need for tools when the model reaches a wrong answer.
Generalization of Tool Learning.
To analyze the generalization ability of Trice, we extend the trained model to unseen datasets and tools. As illustrated in Table 4, we evaluate Vicuna on another two math reasoning datasets (MultiArith (Roy and Roth, 2015) and AddSub (Hosseini et al., 2014)) and one LAMA dataset (SQuAD (Petroni et al., 2019)). Our approach enables continuous optimization of the model’s performance on unseen datasets, with Trice-mix yielding superior results compared to Trice-split. This suggests that Trice equips the model with general tool usage capabilities. Furthermore, we steer the model towards unseen tools by simply modifying the instructions. The performance of Vicuna (Trice-mix) augmented by a retriever on HotpotQA (Yang et al., 2018) advances $\uparrow$6.7than the few-shot manner. Despite the disparities between GPT-3.5 on certain datasets, these findings highlight the promise of multi-tool training based on Trice for facilitating the generalization of tool learning.
4.3 Analyses of Selective Tool Learning
Stage-I is the Foundation of Stable Selective Tool Learning.
Figure 3 showcases the performance of Trice at different training stages, with Vicuna as the representative. It is evident that only trained in stage I (Trice-I), the model acquires efficacious tool usage capabilities, resulting in a substantial performance improvement. Upon further training in stage II (Trice-all), the model experiences additional performance enhancements in both the Split and Mix training settings. However, the results obtained solely from stage II (Trice-II) are unsatisfactory, indicating that the initial tool generation ability bestowed upon the model during stage I is crucial for more stable training.
Stage-II Plays a Pivotal Role in Selective Tool Learning.
To investigate how the models learn to use tools selectively, we analyze the tool usage rate statistics of Vicuna during each training stage in Figure 4. After stage I, we notice that the model’s reliance on tools has significantly deepened on most tasks. This indicates that the model effectively learns the pattern of tool usage in stage I. Still, due to the imbalanced data distribution regarding the presence or absence of tools in the training set, instruct-tuning tends to make the model overly dependent on tools. However, after stage II, the model not only shows performance improvement (see Figure 3) but visibly reduces its dependency on tools, which illustrates that the execution feedback can help mitigate the model’s excessive reliance on tools and alleviate error propagation in tool usage. Moreover, it cannot be ignored that the fluctuation of LAMA differs from others. The decision-making process for invoking the QA model poses challenges, leading to insufficient tool learning during stage I. The improvement in tool usage rate during stage II implies that the execution feedback can help address the issue of inadequate tool learning. The above two phenomena highlight the validity of Trice for selective tool usage.
Figure 5: Case study. We mainly show the responses and predictions of stages I and All.
Case Study.
In Figure 5, we present several cases featuring responses and predictions from different stages. Case 1 suggests that stage II can alleviate the insufficient tool learning in stage I, urging the model to seek assistance from tools for questions it struggles to answer. Though stage I equips the model with a certain level of tool generation capability, it may not excel in making optimal decisions about the tool’s input, as shown in Case 2. Stage II mitigates this limitation and enhances the accuracy of tool use. Case 3 confirms that our proposed method enables the model to use tools judiciously. In Case 4, despite having the same tool invocation in both stages I and II, the model may generate completely opposite answers. This indicates that stage II can further optimize the model’s ability to leverage the return results of tools. However, as shown in Case 5, our model still exhibits certain flaws leading to errors in tool usage. We speculate that this could be attributed to the scale of our backbone models, which generally range from 6-7B, potentially limiting their tool learning ability.
5 Discussion
Knowledge Conflicts.
One particularly challenging issue in tool learning is the problem of knowledge conflicts (Qin et al., 2023a) which may derive from the conflicts between model knowledge and augmented knowledge from tools. LLMs need to have the ability to differentiate knowledge from various sources and discern which ones are valuable, which ones are irrelevant, and even which ones may be harmful. Our approach leverages the feedback loop of trial and error (see Figure 2) to learn when to use tools and when not to. The model learns to recognize situations where relying solely on its intrinsic knowledge may not be sufficient and utilizing tools is more reliable. Similarly, it learns to identify scenarios where its own learned knowledge is capable of solving the problem without the need for extensive tool usage (see Figure 4).
Interactive Learning.
Recent NLP witnesses a rapid advancement in interactive learning (Wang et al., 2023). Collaboration among multi-agents (Lin et al., 2023; Liang et al., 2023) and learning from feedback (Chen et al., 2022; Ichter et al., 2022) are the keys to achieving general embodied intelligence as of now. Our approach is a preliminary endeavor to explore the incorporation of embodied methods into tool learning. By leveraging feedback obtained through interactions between the environment (tools) and multi-agents with varying capabilities, we enable language models to learn more desirable execution strategies (see Figure 5).
6 Conclusion
In this paper, we focus on addressing the challenge of selective utilization of tools by LLMs and propose a two-stage end-to-end training framework dubbed Trice to make LLMs better tool learners with execution feedback. Through comprehensive experiments, we show that our method can achieve better performance compared to GPT-3.5. Further analyses illustrate that Trice can selectively use tools by improving the accuracy of tool usage while enhancing insufficient tool learning and mitigating excessive reliance on tools.
Limitations
In this paper, we focus on addressing the challenge of selective utilization of external tools by LLMs and propose a two-stage end-to-end training framework dubbed Trice to make LLMs better tool learners with execution feedback. Despite our best efforts, there may be still some limitations that remain in this paper.
Method.
Our approach can be applied to any tool-learning scenario, including embodied robotics. However, due to the iterative nature of execution feedback, which relies on continuous trial-and-error, it is typically more suitable for computationally feasible virtual environments, while real-world scenarios often require a significant time investment. In the future, we will explore more scientific and intricate feedback mechanisms to address the limitations above.
Language Models.
Given our limited computational resources, we only conduct experiments on three backbone models with scales of 6-7B. In the future, we may advent on LLMs with different architectures and larger scales.
Tasks and Datasets.
Due to the limited resources, we only experiment on four tasks containing eight datasets. There are also numerous tasks and scenarios that require the utilization of more diverse and complex tools. In the future, we will embark on further research endeavors.
Appendix A Appendix
Name Stage-I Stage-II
lora_r 8 8
lora_alpha 16 16
lora_target_modules q_proj v_proj q_proj v_proj
lora_dropout 0.05 0.05
max_length 2048 2048
batch_size_per_device 48 8
gradient_accumulation_steps 8 32
warmup_steps 0 0
epochs 5 2
lr 1e-4, 3e-4 1e-4, 2e-5
$\alpha$ 0.01, 0.1, 1
Table 5: Hyperparameters to train Chatglm-6B.
Name Stage-I Stage-II
lora_r 8 8
lora_alpha 16 16
lora_target_modules q_proj v_proj q_proj v_proj
lora_dropout 0.05 0.05
max_length 512 512
batch_size_per_device 128 8
gradient_accumulation_steps 8 32
warmup_steps 0 0
epochs 5 2
lr 1e-4, 2e-5 1e-4, 2e-5
$\alpha$ 0.01, 0.1, 1
Table 6: Hyperparameters to train Alpaca-7B and Vicuna-7B.
Figure 6: Training loss variations of Vicuna-7B in stage I of Trice-mix.
Figure 7: Training loss variations of Vicuna-7B in stage II of Trice-mix.
A.1 Task Format
We mainly evaluate our method on four kinds of tasks as shown in Table 2. Eq.1&2 formally define the input and output of each task in general. Here is the detailed format of each task.
Math Reasoning
Instruction $s$: Given a math problem, please solve it and you can use a calculator for help.
Question $q$: Mrs. Hilt has 50 cents. A pencil costs 5 cents. How many pencils can she buy with the money she has?
Tool API $t$ (if needed): calculator(50/5)
Gold Answer $a$: 10
Question Answering
Instruction $s$: Given a question, please answer it and you can use a WikiSearch for help.
Question $q$: Where are sunbeam microwaves made?
Tool API $t$ (if needed): WikiSearch(Sunbeam microwaves manufacturing location)
Gold Answer $a$: Florida
LAMA
Instruction $s$: Given a question, please answer it and you can use a QA model for help.
Question $q$: Winners of the festivals «Chervona Ruta» (Ukraine), «Pearls of the Season» (Ukraine), «Boards» (Moscow), «Woodstock» ( ?
Tool API $t$ (if needed): QA(Where is the Woodstock festival held?)
Gold Answer $a$: Poland
Multilingual QA
Instruction $s$: Given a context, please answer the question in English and you can use a translator for help.
Question $q$: Context: Over the next decade, she went on more than 40 field missions, meeting with refugees and internally displaced persons in over 30 countries. In 2002, when asked what she hoped to accomplish, she stated, “Awareness of the plight of these people. I think they should be commended for what they have survived, not looked down upon.” To that end, her 2001–02 field visits were chronicled in her book Notes from My Travels, which was published in October 2003 in conjunction with the release of her humanitarian drama Beyond Borders. Question: 她在10年内完成了多少任务?
Tool API $t$ (if needed): translator(她在10年内完成了多少任务?)
Gold Answer $a$: more than 40
A.2 Data Preparation
We present the prompt used to generate tool APIs for Math Reasoning, Question Answering, and LAMA in Figure 8-10. Since the sentence to be translated happens to be the provided question, Multilingual QA does not require ChatGPT to generate tool APIs. Due to limited computational resources, we randomly sample train and test sets from each dataset to reduce data scale and training/testing costs. The final data distribution for each task is illustrated in Figure 11.
Figure 8: Prompt used for Math Reasoning to generate tool APIs.
Figure 9: Prompt used for Question Answering to generate tool APIs.
Figure 10: Prompt used for LAMA to generate tool APIs.
Figure 11: Data distribution for each task.
A.3 Response Generation
We show the prompt used to generate candidate responses for ChatGPT and GPT-3.5 in Figure 12-15. We use the same instructions in Figure 5 to generate candidate responses for Vicuna and Alpaca.
Figure 12: Prompt used for Math Reasoning to generate candidate responses.
Figure 13: Prompt used for Question Answering to generate candidate responses.
Figure 14: Prompt used for LAMA to generate candidate responses.
Figure 15: Prompt used for Multilingual QA to generate candidate responses.
A.4 Training Details
The hyperparameters we use to train ChatGLM-6B are shown in Table 5 and Alpaca-7B, Vicuna-7B are shown in Table 6. We present the training loss variations of Vicuna-7B in stages I and II of Trice-mix in Figure 6&7. During training, we observe that despite the decrease in training loss, prolonged reinforcement learning training will result in a significant performance loss. Typically, the model achieves optimal performance within the first 10-40 steps.

### 14 Mar 2024 v3

@@
 To meet this need, we propose Tool leaRning wIth exeCution fEedback (Trice), a two-stage end-to-end framework that enables the model to continually learn through feedback derived from tool execution, thereby learning when and how to use tools effectively.
-Experimental results, backed by further analysis, show that Trice can make the large language model selectively use tools by improving the accuracy of tool usage while enhancing insufficient tool learning and mitigating excessive reliance on tools [^1] Code and datasets are in https://github.com/zjunlp/TRICE..
+Experimental results, backed by further analysis, show that Trice can make the large language model selectively use tools by improving the accuracy of tool usage while enhancing insufficient tool learning and mitigating excessive reliance on tools [^1] Code: https://github.com/zjunlp/TRICE..
 Making Language Models Better Tool Learners with Execution Feedback
-Shuofei ${}^{\spadesuit}$, Honghao ${}^{\spadesuit}$, Chengfei ${}^{\clubsuit}$, Qianghuai ${}^{\clubsuit}$, Huajun ${}^{\spadesuit\heartsuit}$, Ningyu ${}^{\spadesuit}$thanks:     Corresponding Author.
-${}^{\spadesuit}$ Zhejiang University ${}^{\heartsuit}$ Donghai Laboratory ${}^{\clubsuit}$ Alibaba Group {shuofei,guihonghao,huajunsir,zhangningyu}@zju.edu.cn
+Shuofei Qiao1,2, Honghao Gui1,2, Chengfei Lv4, Qianghuai Jia4, Huajun Chen1,2,3, Ningyu Zhang1,2thanks:     Corresponding Author.
+1College of Computer Science and Technology, Zhejiang University 2ZJU-Ant Group Joint Research Center for Knowledge Graphs, Zhejiang University 3ZJU-Hangzhou Global Scientific and Technological Innovation Center, Zhejiang University 4Alibaba Group {shuofei,guihonghao,huajunsir,zhangningyu}@zju.edu.cn
 1 Introduction
@@
 We design particular instructions $s$ tailored to each task.
-In the end, we obtain $\mathcal{D}_{\textrm{tool}}=\{(s,q,t,a)\}_{i=1}^{|\mathcal{D}_{\textrm{tool}}|}$ according with Eq.1&2 containing the tool demand information of the specific LLM as we desire [^2] For more details of data preparation, please refer to A.2..
+In the end, we obtain $\mathcal{D}_{\textrm{tool}}=\{(s,q,t,a)\}_{i=1}^{|\mathcal{D}_{\textrm{tool}}|}$ according with Eq.1&2 containing the tool demand information of the specific LLM as we desire [^2] For more details of data preparation, please refer to Appendix A.2..
 Training.
@@
 4.1 Experimental Settings
-Tasks.
+Tasks and Tools.
 As shown in Table 2, we mainly evaluate our method on four tasks with each task specified to an external tool.
@@
 We simply check for the last number predicted by the model for the math reasoning task and check whether the correct answer is within the first twenty words for other tasks.
+The QA model we use for LAMA is a retrieval-augmented LM fine-tuned on Natural Questions (Kwiatkowski et al., 2019) named Atlas (Izacard et al., 2023).
+We use the 600M parameter NLLB (Costa-jussà et al., 2022) as our machine Translation model for Multilingual QA.
 Candidate Response Generation.
@@
 To differentiate whether or not to use tools among candidate responses, we compel ChatGPT and InstructGPT to utilize tools while allowing Alpaca and Vicuna to make the choice of using tools.
-For ChatGPT and InstructGPT, we prompt them with instructions and few-shot examples, and for Alpaca-7B and Vicuna-7B, we fine-tune them on $\mathcal{D}_{\textrm{tool}}$ with LoRA (Hu et al., 2022) for a few steps in order to equip them with initial abilities for question answering and tool generation [^3] For more details of candidate response generation, please refer to A.3..
+For ChatGPT and InstructGPT, we prompt them with instructions and few-shot examples, and for Alpaca-7B and Vicuna-7B, we fine-tune them on $\mathcal{D}_{\textrm{tool}}$ with LoRA (Hu et al., 2022) for a few steps in order to equip them with initial abilities for question answering and tool generation [^3] For more details of candidate response generation, please refer to Appendix A.3..
 Setting Model Math Reasoning Question Answering LAMA Multilingual QA Avg.
@@
 Setups.
+We fine-tune all our models with LoRA Hu et al.
+(2022) in the format proposed in Alpaca .
 All the models are trained for 5 epochs in stage-I and 2 epochs in stage-II.
@@
 The $\alpha$ is set to {0.01, 0.1, 1} for all the models.
-The detailed hyper-parameters we use are shown in A.4.
+The detailed hyper-parameters we use are shown in Appendix A.4.
 Since sampling responses and training are separated, our whole training procedure only needs to load one model.
@@
 Meanwhile, across all backbone models, Trice-split demonstrates significant improvements compared to the prompt-based baselines, surpassing the few-shot setting with $\uparrow$14.0for ChatGLM, $\uparrow$15.3for Alpaca, and $\uparrow$11.9for Vicuna.
-This indicates that Trice consistently empowers LLMs to use tools effectively, irrespective of the model architecture and scale.
+This indicates that Trice consistently empowers LLMs to use tools effectively, irrespective of the model architecture and scale (ChatGLM-6B is encoder-decoder, while Alpaca-7B and Vicuna-7B are decoder-only).
 Moreover, whether it is completely independent (0% Tool) or dependent (100% Tool) on tools, supervised fine-tuning fails to beat Trice-based training, which highlights the necessity and efficacy of judicious tool learning.
@@
 Table 4: Performance to unseen datasets and tools.
+The presence of an empty value in the Unseen Tool TRICE-APLIT section is due to our reliance on the generalization achieved through mixed-tool training (MIX).
+Testing the new tool individually through separate training (SPLIT) using any single tool would not be suitable.
 Figure 3: Performance of Trice across all tasks at different training stages.
@@
 However, the results obtained solely from stage II (Trice-II) are unsatisfactory, indicating that the initial tool generation ability bestowed upon the model during stage I is crucial for more stable training.
+Figure 5: Case study. We mainly show the responses and predictions of stages I and All.
 Stage-II Plays a Pivotal Role in Selective Tool Learning.
@@
 The above two phenomena highlight the validity of Trice for selective tool usage.
-Figure 5: Case study. We mainly show the responses and predictions of stages I and All.
 Case Study.
@@
 Knowledge Conflicts.
-One particularly challenging issue in tool learning is the problem of knowledge conflicts (Qin et al., 2023a) which may derive from the conflicts between model knowledge and augmented knowledge from tools.
+In tool learning, LLMs manipulate tools and respond to users conditioned on a variety of knowledge sources.
+One particularly challenging issue in tool learning is the problem of knowledge conflicts (Qin et al., 2023a) which may derive from the conflicts between model knowledge and augmented knowledge from tools, and among augmented knowledge from different tools.
+This may lead to a lack of explainability in model prediction and planning.
 LLMs need to have the ability to differentiate knowledge from various sources and discern which ones are valuable, which ones are irrelevant, and even which ones may be harmful.
+This ability becomes even more critical in highly specialized fields such as biomedical research and legal assistance, where the accuracy and reliability of knowledge are of utmost importance.
 Our approach leverages the feedback loop of trial and error (see Figure 2) to learn when to use tools and when not to.
 The model learns to recognize situations where relying solely on its intrinsic knowledge may not be sufficient and utilizing tools is more reliable.
-Similarly, it learns to identify scenarios where its own learned knowledge is capable of solving the problem without the need for extensive tool usage (see Figure 4).
+Similarly, it learns to identify scenarios where its own learned knowledge is capable of solving the problem without the need for extensive tool usage.
+This learning process allows the model to adapt and make informed decisions about when to rely on its own capabilities and when to utilize tools effectively (see Figure 4).
 Interactive Learning.
-Recent NLP witnesses a rapid advancement in interactive learning (Wang et al., 2023).
+Recent NLP has witnessed rapid advancement in interactive learning which considers language models as agents capable of observing, acting, and receiving feedback in a loop with external objects such as humans, knowledge bases, tools, models, and environments (Wang et al., 2023).
 Collaboration among multi-agents (Lin et al., 2023; Liang et al., 2023) and learning from feedback (Chen et al., 2022; Ichter et al., 2022) are the keys to achieving general embodied intelligence as of now.
@@
 By leveraging feedback obtained through interactions between the environment (tools) and multi-agents with varying capabilities, we enable language models to learn more desirable execution strategies (see Figure 5).
+However, our current method is unable to learn the usage of multi-tool compositions.
+In the future, more sophisticated trial-and-error processes and feedback mechanisms will be necessary for LLMs to better utilize tools (e.g.
+learning multi-tool compositions) and even create new tools.
 6 Conclusion
@@
 In the future, we will embark on further research endeavors.
+Acknowledgements
+We would like to express gratitude to the anonymous reviewers for their kind comments.
+This work was supported by the National Natural Science Foundation of China (No.
+62206246), the Fundamental Research Funds for the Central Universities (226-2023-00138), Zhejiang Provincial Natural Science Foundation of China (No.
+LGG22F030011), CAAI-Huawei MindSpore Open Fund, Ningbo Natural Science Foundation (2021J190), Yongjiang Talent Introduction Programme (2021A-156-G), CCF-Tencent Rhino-Bird Open Research Fund, and Information Technology Center and State Key Lab of CAD&CG, Zhejiang University.
 Appendix A Appendix